OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 625 and a3 = 25?

OpenStudy (anonymous):

a1 = 625 a3 = sqrt of a1 a5 = sqrt of a3 a7 = sqrt of a5

Parth (parthkohli):

a5 = 1 a7 = 1/25

Parth (parthkohli):

a6 = 1/5

OpenStudy (anonymous):

a1 = 625 a3 = 625 / q² .... ( q = 5) ....

OpenStudy (anonymous):

a=625 625r^2=25 r^2=5/125 r=sqrt(5/125) ar^6=625(sqrt(5/125)^5) Tap on your calculator and you should get your answer.

OpenStudy (anonymous):

And sorry... the ar^6 should be ar^5. The formula is ar^(n-1) after all.

OpenStudy (anonymous):

so what do i plug in the calculator?

Parth (parthkohli):

\[\LARGE a _{x} = {3125 \over 5x}\]

OpenStudy (anonymous):

yeah and thats the same for ax = 625/x ... you just multiplied 5 :)

OpenStudy (anonymous):

625 isnt right

OpenStudy (anonymous):

Key in this: 625(sqrt(5/125)^5) sqrt is the square root.